OpenStudy (anonymous):

The area of a rectangular swimming pool is 10x2 – 19x – 15. The length of the pool is 5x + 3. What is the width of the pool? 2x – 18 2x – 5 5x – 5 5x – 22

OpenStudy (anonymous):

Area is length times width. So if length is 5x+3 you have to multiply that by something that gives you 10x^2-19x-15. Can you figure it out now?

OpenStudy (anonymous):

no, i dont understand this stuff.

OpenStudy (anonymous):

You have to multiply 5x+3 by one of those possible answers to get the area, which is 10x^2-19x-15. Do you know how to use the FOIL method? Try multiplying 5x+3 by 2x-18 and 2x-5. One of those is the answer.

OpenStudy (anonymous):

ive heard of the foil method but im not sure how to use it.

OpenStudy (anonymous):

Ok, let's do 5x+3 * 2x-5. Foil stands for First, Outside, Inside, Last. First term times first term is 5x*2x = 10x^2. Outside terms: 5x*-5= -25x Inside terms 3*2x = 6x Last terms 3*-5=-15 whole thing is 10x^2-25x+6x-15 which simplifies to the given area. So the answer is 2x-5. Understand?

OpenStudy (anonymous):

oh yeah, you have to do it that way. thanks i remember the foil method now!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Keep it up!